OpenStudy (anonymous):

plzzzz help z[8.01, 8.07, 10.01]Which of the following represents x squared minus sixteen all over x squared minus four x in simplified form?

OpenStudy (bibby):

\(a^2-b^2=(a-b)*(a+b)\\\huge \frac{x^2-16}{x^2-4}\) theyre both in that form

OpenStudy (bibby):

hmm?

OpenStudy (bibby):

the question marks disappear if you refresh but idk what you were doing

OpenStudy (bibby):

anyhow my point was, we can rewrite \(\huge \frac{x^2-16}{x^2-4}\) in \(a^2-b^2\) form let's start with \(x^2-16\)

OpenStudy (anonymous):

\[\frac{ x^2{-16} }{ x^{2}-4x}\]

OpenStudy (anonymous):

thats how it looks

OpenStudy (bibby):

oh whatever, no big deal

OpenStudy (bibby):

\(x^2-16=(x)^2-(?)^2\) what goes in t he question mark?

OpenStudy (anonymous):

4?

OpenStudy (bibby):

exactly. so now we have \(a^2-b^2=(a-b)*(a+b)\) and \((x^2-4^2)\) factorize the numerator

OpenStudy (anonymous):

hu?

OpenStudy (bibby):